# Nearly-polynomial inverse theorem for the $U^{d}$ norm in degree $d+1$

### 17 Mar 2026 v1

@@ 4.2 Lower-degree correlation
Recall that standard correlation results show that if a polynomial $f$ has a short decomposition in terms of lower degree polynomials, then it has a significant correlation with a polynomial of degree at most $\deg(f)-1$. In this section we show that if a polynomial $f$ has a short $\operatorname{rk}^{*}$-decomposition, then it has a significant correlation with a polynomial of degree at most $\deg(f)-2$. This relies on our correlation lemma in Lemma 3.1, but here we only assume small $\operatorname{rk}^{*}$.
@@
In order to convert a short $\operatorname{rk}^{*}$-decomposition to significant correlation, via Lemma 3.1, we use the following lemma which gives an essentially optimal lower bound on the joint zero probability $\mathbb{P}(\operatorname{Z}(\mathbf{A}))$.
@@
Put $r=\operatorname{rk}^{*}(P)$ and $\mathbb{F}=\mathbb{F}_{q}$. Write $f=\sum_{i=1}^{r}\alpha_{i}\beta_{i}$ with $1\leq\deg(\alpha_{i})\leq\deg(\beta_{i})\leq k-1$ and $\deg(\alpha_{i}\beta_{i})\leq k$. Assume without loss of generality that all polynomials $\alpha_{i}$ satisfying both $\deg(\alpha_{i})=1$ and $\deg(\alpha_{i}\beta_{i})=k$ are precisely $\alpha_{t+1},\ldots,\alpha_{r}$ for some $t\leq r$. Observe that for every $i\leq t$ we necessarily have $\deg(\beta_{i})=\deg(\alpha_{i}\beta_{i})-\deg(\alpha_{i})\leq k-2$.
Put $\mathbf{A}^{1}=(\alpha_{1},\ldots,\alpha_{t})$, $\mathbf{A}^{2}=(\alpha_{t+1},\ldots,\alpha_{r})$, and $\mathbf{B}^{1}=(\beta_{1},\ldots,\beta_{t})$. Since
@@
$\mathbb{P}\big((\mathbf{A}^{1},\mathbf{B}^{1})=(y^{1},z^{1})\,|\,\mathbf{A}^{2}=0\big)\geq q^{-2t}.$ (4)
Put $y=(y^{1},y^{2})\in\mathbb{F}^{r}$ with $y^{2}=0\in\mathbb{F}^{r-t}$. We decompose $f$ as follows (recall $y_{i}=0$ for $i\geq t$):
@@
Let $A_{i}=\alpha_{i}-y_{i}$ for $i\leq r$, let $A_{r+i}=\beta_{i}-z_{i}$ for $i\leq t$, and let $c=\sum_{i=1}^{t}y_{i}z_{i}$ ($\in\mathbb{F}$). Then $f-c\in\langle A_{1},...,A_{m}\rangle$ with $m=r+t\leq 2r$. Moreover, for $i\leq k$ we have $\deg(A_{r+i})=\deg(\beta_{i})\leq k-2$, and for $i\leq r$ we have $\deg(A_{i})=\deg(\alpha_{i})\leq\lfloor k/2\rfloor\leq k-2$, using our assumption $k\geq 3$.
Finally, by the definition of $\operatorname{rk}^{*}(f)$, the linear components of $\alpha_{t+1},\ldots,\alpha_{r}$ are linearly independent. Thus, the affine map $\mathbf{A}^{2}\colon\mathbb{F}^{n}\to\mathbb{F}^{r-t}$ is perfectly equidistributed, and in particular, $\mathbb{P}(\mathbf{A}^{2}=0)\geq q^{-(r-t)}$. Therefore,
$\mathbb{P}(\forall i\colon A_{i}=0)$ $=\mathbb{P}((\mathbf{A}^{1},\mathbf{A}^{2},\mathbf{B}^{1})=(y^{1},y^{2},z^{1}))$
$=\mathbb{P}(\mathbf{A}^{2}=0)\mathbb{P}\big((\mathbf{A}^{1},\mathbf{B}^{1})=(y^{1},z^{1})\,|\,\mathbf{A}^{2}=0\big)\geq q^{-(r-t)}q^{-2t}=q^{-m},$
where the inequality combines the equidistribution of $\mathbf{A}^{2}$ together with (4). This completes the proof. ∎
@@
Since $\sum_{i=1}^{m}c_{i}A_{i}\in\operatorname{Poly}_{k-2}$, we deduce that
@@ 5.1 Discrete derivatives: Proofs
$\Delta^{d}_{\mathbf{x}}f(x^{(d+1)})=\sum_{I}c_{I}\sum_{\varphi\colon[k]\twoheadrightarrow[d]}\,\prod_{j=1}^{k}x_{I_{j}}^{(\varphi(j))}.$
@@
$\Delta^{d}_{\mathbf{x}}f(x^{(d+1)})=\sum_{\varphi\colon[k]\twoheadrightarrow[d]}\,\prod_{j=1}^{k}x_{i_{j}}^{(\varphi(j))}.$
@@
For the induction step, we have
$\Delta^{d}_{\mathbf{x}}f(x^{(d+1)})$ $=\Delta_{x^{(d)}}\Delta^{d-1}_{\mathbf{x}}f(x^{(d+1)})=\Delta^{d-1}_{\mathbf{x}}f(x^{(d+1)}+x^{(d)})-\Delta^{d-1}_{\mathbf{x}}f(x^{(d+1)})$
@@
where the third equality uses the induction hypothesis, and the last equality uses the natural bijection between semi-surjections $\varphi^{\prime}\colon[k]\twoheadrightarrow[d]$ and pairs of semi-surjections $(\varphi,\tau)$ where $\varphi\colon[k]\twoheadrightarrow[d-1]$ and $\tau\colon\varphi^{\raisebox{0.60275pt}{$\scriptscriptstyle\!-\!1$}}(d)\to\{d,d+1\}$ with $d\in\operatorname{Im}(\tau)$. This completes the induction step and the proof.
@@
$\overline{g}(x^{(1)},\ldots,x^{(k)})=\sum_{\pi\in S_{k}}\,\prod_{j=1}^{k}x_{i_{j}}^{(\pi(j))}\,.$
@@
It suffices to show $\partial_{x_{n}^{(k)}}\overline{g}(\mathbf{x})=\overline{\partial_{x_{n}}g}(\mathbf{x})$ for $g(x)=\prod_{j=1}^{k}x_{i_{j}}$ ($i_{1}\leq\cdots\leq i_{k}$) a single monomial, and $\mathbf{x}=(x^{(1)},\ldots,x^{(k)})$, since the result would follow by symmetry and linearity. On the one hand, [^9] For example, $\partial_{x_{2}}(x_{1}x_{2}^{2})=x_{1}x_{2}+x_{1}x_{2}$.
@@
is multilinear, so its only monomials containing $x^{(k)}_{n}$ are those where an occurrence of $x_{n}$ is replaced by $x_{n}^{(k)}$; therefore,
$\partial_{x^{(k)}_{n}}\overline{g}(x^{(1)},\ldots,x^{(k)})=\sum_{\begin{subarray}{c}t\in[k]\\ i_{t}=n\end{subarray}}\,\sum_{\begin{subarray}{c}\sigma\in S_{k}\\ \pi(t)=k\end{subarray}}\,\prod_{\begin{subarray}{c}j=1\\ j\neq t\end{subarray}}^{k}x_{i_{j}}^{(\sigma(j))}=\sum_{\begin{subarray}{c}t\in[k]\\ i_{t}=n\end{subarray}}\,\sum_{\begin{subarray}{c}\sigma\in S_{k}\\ \pi(t)=k\end{subarray}}\,\prod_{j=1}^{k-1}x_{i_{j}}^{(\sigma(j))}\;.$ (6)
@@
By linearity, we may assume that $f$ is a single monomial, $f=\prod_{j=1}^{d+1}x_{i_{j}}$, of degree $d+1$. For a function $\varphi\colon[k]\to[d+1]$, we abbreviate $f^{(\varphi)}=\prod_{j=1}^{d+1}x_{i_{j}}^{(\varphi(j))}$. By Lemma 5.5,
@@
where $S_{d}^{t}$ denotes the set of all surjections $\varphi\colon[d+1]\to[d]$ satisfying $|\varphi^{\raisebox{0.60275pt}{$\scriptscriptstyle\!-\!1$}}(t)|=2$. We claim that for any $t\in[d]$, there is a 22-to-11 map from $S_{d+1}$ to $S_{d}^{t}$. Indeed, we map each permutation $\sigma\in S_{d+1}$ to the function $\sigma_{t}\in S_{d}^{t}$ given by
@@
Put differently, $\sigma_{t}(x)=\sigma_{\sigma^{\raisebox{0.43054pt}{$\scriptscriptstyle\!-\!1$}}(d+1)\to t}$, where for any function $h$ we write $h_{x_{0}\to y_{0}}$ for
$h_{x_{0}\to y_{0}}(x)=\begin{cases}y_{0}&\text{if }x=x_{0}\\ h(x)&\text{else}\end{cases}\;.$
To see that this map is 22-to-11, let $\varphi\in S_{d}^{t}$, and consider the two functions $\varphi_{a\mapsto d+1}$ and $\varphi_{b\mapsto d+1}$. That both functions are in $S_{d+1}$ follows from our assumption on $\varphi$. A moment’s thought reveals that both—and none others—are mapped to $\varphi$, that is, $\sigma_{\sigma^{\raisebox{0.43054pt}{$\scriptscriptstyle\!-\!1$}}(d+1)\to t}=\varphi$ if and only if $\sigma=\varphi_{a\mapsto d+1}$ or $\sigma=\varphi_{b\mapsto d+1}$, proving our claim. Going back to $\Delta^{d}_{\mathbf{x}}f(x^{(d+1)})$, we thus have
@@
using the multilinearity of $\overline{f}$ in the last equality. This completes the proof. ∎
@@ 6.1 Bias of general polynomials
While the bias of a polynomial is closely related to the bias of its top-degree component, it generally has little to do with the bias of its lower-degree components. For example, consider the univariate polynomial $x^{2}+x$, over $\mathbb{F}_{p}$ with $p>2$. We have
$\operatorname{bias}(x^{2}+x)\cdot p=\Big|\sum_{x\in\mathbb{F}_{p}}\omega_{p}^{x^{2}+x}\Big|=\Big|\omega_{p}^{-1/4}\sum_{x\in\mathbb{F}_{p}}\omega_{p}^{(x+1/2)^{2}}\Big|=\Big|\sum_{y\in\mathbb{F}_{p}}\omega_{p}^{y^{2}}\Big|=\sqrt{p},$
where the last step uses the formula for a basic quadratic Gauss sum. Thus, $\operatorname{bias}(x^{2}+x)=\operatorname{bias}(x^{2})=p^{-1/2}$, whereas for the lower degree component of $x^{2}+x$ we have $\operatorname{bias}(x)=0$.
@@
Observe that for any fixed $y\in E^{c}$, $A(x,y)+B(y)$ is polynomial in $x$ of degree exactly $1$, and thus takes any value in the field with the same probability. Furthermore observe that
$\operatorname{bias}(A)=\operatorname*{\mathbb{E}}(\chi(A))=\operatorname*{\mathbb{E}}_{y}\operatorname*{\mathbb{E}}_{x}(\chi(A(x,y)))=\operatorname*{\mathbb{E}}_{y}\mathbf{1}_{E}(y)=\mathbb{P}(E).$
@@
$\operatorname{bias}(A+B)$ $=\operatorname*{\mathbb{E}}(\chi(A+B))=\mathbb{P}(E)\operatorname*{\mathbb{E}}(\chi(A+B)\,|\,E)+\mathbb{P}(E^{c})\operatorname*{\mathbb{E}}(\chi(A+B)\,|\,E^{c})$
$=\operatorname{bias}(A)\operatorname{bias}(B\,|\,E)+\mathbb{P}(E^{c})\cdot 0.\qed$
Our main result in this subsection bounds the bias of the $d$-times iterated derivative of $f\in\operatorname{Poly}_{d+1}$ in terms of the components of $f$.
@@
For $f\in\operatorname{Poly}_{d+1}(\mathbb{F})$ where $\operatorname{ch}(\mathbb{F})\neq 2$, with $g=f_{d+1}$ and $h=f_{d}$,
@@
$\operatorname{\mathfrak{bias}}(\Delta^{d}_{\mathbf{v}}f(x))=\operatorname*{\mathbb{E}}_{x}\operatorname*{\mathbb{E}}_{\mathbf{v}}\chi(\overline{g}(\mathbf{v},x^{\prime})+\overline{h}(\mathbf{v}))=\operatorname*{\mathbb{E}}_{x^{\prime}}\operatorname*{\mathbb{E}}_{\mathbf{v}}\chi(\overline{g}(\mathbf{v},x^{\prime})+\overline{h}(\mathbf{v}))=\operatorname{\mathfrak{bias}}(\overline{g}+\overline{h}).$
@@
These complete the proof. ∎

### 30 Apr 2026 v2

@@ 4.2 Lower-degree correlation
Recall that standard correlation results show that if a polynomial $f$ has a short decomposition in terms of lower degree polynomials, then it has a significant correlation with a polynomial of degree at most $\deg(f)-1$. In this section we show that if a polynomial $f$ has a short $\operatorname{rk}^{*}$-decomposition, then it has a significant correlation with a polynomial of degree at most $\deg(f)-2$. [^22] Assuming $\deg(f)\geq 3$. If $\deg(f)=2$ we only get correlation with degree $\deg(f)-1$, naturally. This relies on our correlation lemma in Lemma 3.1, but here we only assume small $\operatorname{rk}^{*}$.
@@
For completeness, if $\operatorname{rk}^{*}(f)=\infty$ then the right hand side should be interpreted as 0.
In order to convert a short $\operatorname{rk}^{*}$-decomposition to significant correlation, via Lemma 3.1, we use the following lemma which puts a shift of $f$ in an ideal generated by few polynomials of degree at most $\deg(f)-2$ that, moreover, frequently vanish together.
@@
Put $r=\operatorname{rk}^{*}(f)$ and $\mathbb{F}=\mathbb{F}_{q}$. Assume $r<\infty$, as otherwise we are done by taking $\mathbf{A}=(x_{1},\ldots,x_{n})$ and $c=f(0)$. Assume $f$ is non-constant, as otherwise we are done by taking $m=0$ and $c=f\in\mathbb{F}_{q}$. Write $f=\sum_{i=1}^{r}\alpha_{i}\beta_{i}$ with $1\leq\deg(\alpha_{i})\leq\deg(\beta_{i})\leq k-1$ and $\deg(\alpha_{i}\beta_{i})\leq k$. Assume without loss of generality that the polynomials $\alpha_{i}$ satisfying both $\deg(\alpha_{i})=1$ and $\deg(\alpha_{i}\beta_{i})=k$ are precisely $\alpha_{t+1},\ldots,\alpha_{r}$ for some $t\leq r$. Observe that for every $i\leq t$ we necessarily have $\deg(\beta_{i})=\deg(\alpha_{i}\beta_{i})-\deg(\alpha_{i})\leq k-2$; indeed, we have that $\deg(\alpha_{i}\beta_{i})\leq k-1$ or $\deg(\alpha_{i})\geq 2$.
Let $\mathbf{A}^{1}=(\alpha_{1},\ldots,\alpha_{t})$, $\mathbf{A}^{2}=(\alpha_{t+1},\ldots,\alpha_{r})$, and $\mathbf{B}^{1}=(\beta_{1},\ldots,\beta_{t})$, and $m=r+t$. We claim that there are $y^{1},z^{1}\in\mathbb{F}^{t}$ such that the map $(\mathbf{A}^{1},\mathbf{A}^{2},\mathbf{B}^{1})\colon\mathbb{F}^{n}\to\mathbb{F}^{m}$ satisfies
$\mathbb{P}\big((\mathbf{A}^{1},\mathbf{A}^{2},\mathbf{B}^{1})=(y^{1},0,z^{1})\big)\geq q^{-m}.$ (4)
If $t=r$ (for example, when $\deg(f)<k$), that is, $\mathbf{A}^{2}$ is empty, then (4) follows by the pigeonhole principle, so henceforth assume otherwise. By the definition of $\operatorname{rk}^{*}(f)$, the linear components of $\alpha_{t+1},\ldots,\alpha_{r}$ are linearly independent, which implies that the affine map $\mathbf{A}^{2}\colon\mathbb{F}^{n}\to\mathbb{F}^{r-t}$ is perfectly equidistributed; indeed, the associated linear map is surjective, so its fibers are all cosets of the kernel and thus of the same size. In particular, $\mathbb{P}(\mathbf{A}^{2}=0)=q^{-(r-t)}$ (>0>0). Since
@@
$\mathbb{P}\big((\mathbf{A}^{1},\mathbf{B}^{1})=(y^{1},z^{1})\,|\,\mathbf{A}^{2}=0\big)\geq q^{-2t}.$
This proves (4):
$\mathbb{P}\big((\mathbf{A}^{1},\mathbf{A}^{2},\mathbf{B}^{1})=(y^{1},0,z^{1})\big)$ $=\mathbb{P}(\mathbf{A}^{2}=0)\cdot\mathbb{P}\big((\mathbf{A}^{1},\mathbf{B}^{1})=(y^{1},z^{1})\,|\,\mathbf{A}^{2}=0\big)$
$\geq q^{-(r-t)}q^{-2t}=q^{-m}.$
Now, put $y=(y^{1},y^{2})\in\mathbb{F}^{r}$ with $y^{2}=0\in\mathbb{F}^{r-t}$. Decompose $f$ as follows (recall $y_{i}=0$ for $i>t$):
@@
Let $A_{i}=\alpha_{i}-y_{i}$ for $i\leq r$, let $A_{r+i}=\beta_{i}-z_{i}$ for $i\leq t$, and let $c=\sum_{i=1}^{t}y_{i}z_{i}$ ($\in\mathbb{F}$). Then $f-c\in\langle A_{1},...,A_{m}\rangle$ with $m=r+t\leq 2r$. Moreover, for $i\leq t$ we have $\deg(A_{r+i})=\deg(\beta_{i})$, and $1\leq\deg(\beta_{i})\leq k-2$, while for $i\leq r$ we have $\deg(A_{i})=\deg(\alpha_{i})$, and $1\leq\deg(\alpha_{i})\leq\lfloor k/2\rfloor\leq k-2$ using our assumption $k\geq 3$. Observe that
$\mathbb{P}(\forall i\colon A_{i}=0)=\mathbb{P}\big((\mathbf{A}^{1},\mathbf{A}^{2},\mathbf{B}^{1})=(y^{1},y^{2},z^{1})\big)\geq q^{-m},$
where the inequality is (4). This completes the proof. ∎
@@
Since $\sum_{i=1}^{m}c_{i}A_{i}\in\operatorname{Poly}_{k-2}$, we deduce that $f$ is significantly correlated with a polynomial of degree at most k2k-2:
@@ 5.1 Discrete derivatives: Proofs
$\Delta^{d}_{x^{(1)},\ldots,x^{(d)}}f(x^{(d+1)})=\sum_{I}c_{I}\sum_{\varphi\colon[k]\twoheadrightarrow[d]}\,\prod_{j=1}^{k}x_{I_{j}}^{(\varphi(j))}.$
@@
$\Delta^{d}_{x^{(1)},\ldots,x^{(d)}}f(x^{(d+1)})=\sum_{\varphi\colon[k]\twoheadrightarrow[d]}\,\prod_{j=1}^{k}x_{i_{j}}^{(\varphi(j))}.$
@@
For the induction step, abbreviate $\mathbf{x}=(x^{(1)},\ldots,x^{(d-1)})$, and observe that
$\Delta^{d}_{x^{(1)},\ldots,x^{(d)}}f(x^{(d+1)})$ $=\Delta_{x^{(d)}}\Delta^{d-1}_{\mathbf{x}}f(x^{(d+1)})=\Delta^{d-1}_{\mathbf{x}}f(x^{(d+1)}+x^{(d)})-\Delta^{d-1}_{\mathbf{x}}f(x^{(d+1)})$
@@
where the third equality uses the induction hypothesis twice, and the last equality uses the natural bijection between semi-surjections $\varphi^{\prime}\colon[k]\twoheadrightarrow[d]$ and pairs $(\varphi,\tau)$ where $\varphi\colon[k]\twoheadrightarrow[d-1]$ and $\tau\colon\varphi^{\raisebox{0.60275pt}{$\scriptscriptstyle\!-\!1$}}(d)\to\{d,d+1\}$ with $d\in\operatorname{Im}(\tau)$. This completes the induction step and the proof.
@@
$\overline{g}(x^{(1)},\ldots,x^{(k)})=\sum_{\sigma\in S_{k}}\,\prod_{j=1}^{k}x_{i_{j}}^{(\sigma(j))}\,.$
@@
It suffices to show $\partial_{x_{n}^{(k)}}\overline{g}(\mathbf{x})=\overline{\partial_{x_{n}}g}(\mathbf{x})$ for $g(x)=\prod_{j=1}^{k}x_{i_{j}}$ ($i_{1}\leq\cdots\leq i_{k}$) a single monomial, and $\mathbf{x}=(x^{(1)},\ldots,x^{(k-1)})$, since the result would follow by symmetry and linearity (the polarization $\overline{g}$ is symmetric by Fact 5.6; directional derivatives are linear in $c$). On the one hand, [^23] For example, $\partial_{x_{2}}(x_{1}x_{2}^{2})=x_{1}x_{2}+x_{1}x_{2}$.
@@
is multilinear, so its only monomials containing $x^{(k)}_{n}$ are those where an occurrence of $x_{n}$ is replaced by $x_{n}^{(k)}$ (that is, $i_{t}=n$ and $\sigma(t)=k$); therefore,
$\partial_{x^{(k)}_{n}}\overline{g}(x^{(1)},\ldots,x^{(k-1)})=\sum_{\begin{subarray}{c}t\in[k]\\ i_{t}=n\end{subarray}}\,\sum_{\begin{subarray}{c}\sigma\in S_{k}\\ \sigma(t)=k\end{subarray}}\,\prod_{\begin{subarray}{c}j=1\\ j\neq t\end{subarray}}^{k}x_{i_{j}}^{(\sigma(j))}=\sum_{\begin{subarray}{c}t\in[k]\\ i_{t}=n\end{subarray}}\,\sum_{\begin{subarray}{c}\sigma\in S_{k}\\ \sigma(t)=k\end{subarray}}\,\prod_{j=1}^{k-1}x_{i_{j}}^{(\sigma(j))}\;.$ (6)
@@
By linearity, we may assume that we are given a single monomial, $f=\prod_{j=1}^{d+1}x_{i_{j}}$, of degree $d+1$. We need to show, for $\mathbf{x}=(x^{(1)},\ldots,x^{(d)})$, that
$\Delta^{d}_{\mathbf{x}}f(x^{(d+1)})=\overline{f}\Big(\mathbf{x},\,x^{(d+1)}+\frac{1}{2}\sum_{t=1}^{d}x^{(t)}\Big).$
For a function $\varphi\colon[d+1]\to[d+1]$, we abbreviate $f^{(\varphi)}=\prod_{j=1}^{d+1}x_{i_{j}}^{(\varphi(j))}$. Observe that every semi-surjection $\varphi\colon[d+1]\twoheadrightarrow[d]$ is either a permutation or takes exactly one value twice. Therefore, by Lemma 5.5,
@@
where $S_{d}^{t}$ denotes the set of all surjections $\varphi\colon[d+1]\to[d]$ satisfying $|\varphi^{\raisebox{0.60275pt}{$\scriptscriptstyle\!-\!1$}}(t)|=2$. For $t\in[d]$ and a permutation $\sigma\in S_{d+1}$, consider the function $\sigma_{t}\in S_{d}^{t}$ given by
@@
We claim that the map $\sigma\mapsto\sigma_{t}$, from $S_{d+1}$ to $S_{d}^{t}$, is 22-to-11. Indeed, for any $\varphi\in S_{d}^{t}$, write $\varphi^{-1}(t)=\{a,b\}$ and consider the two permutations
$\varphi_{a\mapsto d+1}(x)=\begin{cases}d+1&\text{if }x=a\\ \varphi(x)&\text{else}\end{cases}\quad\text{ and }\quad\varphi_{b\mapsto d+1}(x)=\begin{cases}d+1&\text{if }x=b\\ \varphi(x)&\text{else}\end{cases}.$
A moment’s thought reveals that $\sigma\in S_{d+1}$ satisfies $\sigma_{t}=\varphi$ if and only if $\sigma\in\{\varphi_{a\mapsto d+1},\varphi_{b\mapsto d+1}\}$.
Going back to $\Delta^{d}_{\mathbf{x}}f(x^{(d+1)})$, and using the fact that $\operatorname{ch}(\mathbb{F})\neq 2$, we thus have
$\sum_{\varphi\in S_{d}^{t}}\,f^{(\varphi)}=\frac{1}{2}\sum_{\sigma\in S_{d+1}}f^{(\sigma_{t})}.$
Therefore,
@@
using Fact 5.6 in the second equality, and the multilinearity of $\overline{f}$ in the last equality. This completes the proof. ∎
@@ 6.1 Bias of general polynomials
While the bias of a polynomial is closely related to the bias of its top-degree component, it can have little to do with the bias of its lower-degree components. For example, consider the univariate polynomial $x^{2}+x$, over $\mathbb{F}_{p}$ with $p>2$. We have
$\operatorname{bias}(x^{2}+x)\cdot p=\Big|\sum_{x\in\mathbb{F}_{p}}e(x^{2}+x)\Big|=\Big|e(-1/4)\sum_{x\in\mathbb{F}_{p}}e((x+1/2)^{2})\Big|=\Big|\sum_{y\in\mathbb{F}_{p}}e(y^{2})\Big|=\sqrt{p},$
where the last step uses the formula for a basic quadratic Gauss sum. Thus, $\operatorname{bias}(x^{2}+x)=\operatorname{bias}(x^{2})=p^{-1/2}$. However, for the lower degree component of $x^{2}+x$ we have $\operatorname{bias}(x)=0$.
@@
Observe that for any fixed $y\in E^{c}$, $A(x,y)+B(y)$ has degree exactly $1$ (in $x$), and thus takes any value in the field with the same probability. Furthermore observe that
$\operatorname{\mathfrak{bias}}(A)=\operatorname*{\mathbb{E}}(\chi(A))=\operatorname*{\mathbb{E}}_{y}\operatorname*{\mathbb{E}}_{x}(\chi(A(x,y)))=\operatorname*{\mathbb{E}}_{y}\mathbf{1}_{E}(y)=\mathbb{P}(E).$
@@
$\operatorname{\mathfrak{bias}}(A+B)$ $=\operatorname*{\mathbb{E}}(\chi(A+B))=\mathbb{P}(E)\operatorname*{\mathbb{E}}(\chi(A+B)\,|\,E)+\mathbb{P}(E^{c})\operatorname*{\mathbb{E}}(\chi(A+B)\,|\,E^{c})$
$=\operatorname{\mathfrak{bias}}(A)\operatorname{\mathfrak{bias}}(B\,|\,E)+\mathbb{P}(E^{c})\cdot 0.\qed$
Using Proposition 5.1 and Lemma 6.3, our main result in this subsection bounds the bias of the $d$-times iterated derivative of $f\in\operatorname{Poly}_{d+1}$ in terms of (polarizations of) the components of $f$.
@@
For $f\in\operatorname{Poly}_{d+1}(\mathbb{F})$ where $\operatorname{ch}(\mathbb{F})>2$, with $g=f_{d+1}$ and $h=f_{d}$,
@@
$\operatorname{\mathfrak{bias}}(\Delta^{d}f)=\operatorname{\mathfrak{bias}}(\Delta^{d}_{\mathbf{v}}f(x))=\operatorname*{\mathbb{E}}_{x}\operatorname*{\mathbb{E}}_{\mathbf{v}}\chi(\overline{g}(\mathbf{v},x^{\prime})+\overline{h}(\mathbf{v}))=\operatorname*{\mathbb{E}}_{x^{\prime}}\operatorname*{\mathbb{E}}_{\mathbf{v}}\chi(\overline{g}(\mathbf{v},x^{\prime})+\overline{h}(\mathbf{v}))=\operatorname{\mathfrak{bias}}(\overline{g}+\overline{h}).$
@@
Taking absolute value and using the triangle inequality completes the proof. ∎